# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ 3.2 Bounded Dense Noise
Thus, we obtain $|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|\leq 2m\tau_{0}$ and $\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}\leq 2\sqrt{m}\tau_{0}$, and hence by (3.13) with small enough $c_{4}$ we obtain $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}\leq 4\tau_{0}$. Therefore, Proposition 2.1 implies $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 28\tau_{0}$. By using an argument analogous to (3.8), we arrive at $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 57\tau_{0}$. ∎
The next theorem provides converse bounds that indicate the sharpness of the $O(\tau_{0})$ bounds in Theorems 3.33.4. Using the complex Gaussian $\bm{\Phi}$ and observations $\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$ or $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\tau}$, we show that no algorithm can reconstruct $\mathbf{x}\in\Sigma^{n,*}_{s}$ to an error substantially smaller than $O(\tau_{0})$. The idea is to identify another signal $\mathbf{x}^{\prime}\in\Sigma^{n,*}_{s}$ that is indistinguishable from $\mathbf{x}$ and satisfies $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}=\tilde{\Omega}(\tau_{0})$. We leave the removal of log factors to future work.
@@
On the event $\{|\mathcal{J}_{\mathbf{x},s/4m}|\leq\frac{s}{4}\}$, (3.19) translates into $\frac{s}{2}$ real linear equations, so such $\tilde{\bm{\delta}}$ exists when $s\geq 4$. Now we consider $\mathbf{x}^{\prime}=\frac{\mathbf{x}+\tilde{\bm{\delta}}}{\|\mathbf{x}+\tilde{\bm{\delta}}\|_{2}}\in\Sigma^{n,*}_{s}$, and similarly to (3.18) we have $\frac{\tau_{*}}{2}\leq\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\leq\frac{3\tau_{*}}{2}$, and $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ holds with probability at least $1-\frac{4}{m}$. We further bound $\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}$. For $i\in\mathcal{J}_{\mathbf{x},s/4m}$, we have $\bm{\Phi}_{i}^{*}\tilde{\bm{\delta}}=0$ and hence $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime})=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$. For $i\notin\mathcal{J}_{\mathbf{x},s/4m}$, (2.11) gives
@@
where the last inequality we use $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ and substitute $m=C_{0}s\log(\frac{en}{s})$. It follows that
@@
Simulation: [^6] The MATLAB codes for generating the figures in this paper are available in https://junrenchen58.github.io/. We pause to use experimental results to provide evidence of the achievability and tightness of $O(\tau_{0})$. In all of our experiments, the data points are averaged over $50$ independent trials, each of which concerns the recovery of $\mathbf{x}$ uniformly drawn from $\Sigma^{500,*}_{5}$ from $300$ phase-only measurements. We provide the optimally tuned $\varepsilon$ to basis pursuit (1.8), namely $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$. For the post-sign noise, we test $\tau_{0}\in\{0.04,0.08,0.12,0.16,\cdots,0.36,0.40\}$ and adopt such corruption pattern: find $\theta_{0}\in[0,\frac{\pi}{2}]$ such that $|e^{\mathbf{i}\theta_{0}}-1|=\tau_{0}$ and then corrupt $\mathbf{z}$ to $\breve{\mathbf{z}}=e^{\mathbf{i}\theta_{0}}\mathbf{z}$. For the pre-sign noise, we test $\tau_{0}\in\{0.04,0.12,0.20,0.28,\cdots,0.76,0.84\}$ and generate the noisy observations through $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\tau_{0}\operatorname{sign}(\bm{\Phi}\mathbf{x})\mathbf{i})$. The results are given in Figures 12 and are consistent with our theorems.
@@ 3.3 Sparse Phase Corruption
Consider PO-CS of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\zeta}$ with $\bm{\zeta}$ obeying $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ for some small enough $\zeta_{0}$ and $\|\bm{\zeta}\|_{\infty}\leq 2$. If $m\geq C_{1}s\log(\frac{en}{s})$ with sufficiently large $C_{1}$, then the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, with $\hat{\mathbf{x}}$ being solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};11\zeta_{0}\log(\frac{e}{\zeta_{0}}))$ in (1.8), satisfies
@@
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: The first step is to show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$. As we can use Corollary 1 to show $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},c_{1})$ for sufficiently small $c_{1}$ with the promised probability, we only need to ensure small enough $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$. In the proof of Theorem 3.4 we show (3.12) for $\tilde{\bm{\tau}}_{2}$ satisfying $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$ and $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$. An identical argument yields
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\zeta_{0}\log(e/\zeta_{0})}\right)$
with the probability we promise. Thus, $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ is small enough by the scaling of $m$ and $\zeta_{0}$ being small enough.
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: The second step is to bound the $\ell_{2}$ measurement error at $\mathbf{x}^{\star}=\frac{\kappa m\cdot\mathbf{x}}{\|\bm{\Phi}\mathbf{x}\|_{1}}$,
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq(1+c_{2})\left[\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]$ (3.21)
where $c_{2}>0$ is small enough due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$. We let $\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\in\{-1,1\}^{m}$ whose $1$’s indicate the support set of $\bm{\zeta}$. Then we have
@@
Without loss of generality, we assume $\zeta_{0}m$ is a positive integer. For fixed $I$ with cardinality $\zeta_{0}m$, $\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}$ follows Chi-squared distribution with $2\zeta_{0}m$ degrees of freedom. Then by a standard concentration bound [30, Lem. 1], we obtain that for any $t\geq 0$ and any $I\subset[m]$ with $|I|=\zeta_{0}m$,
@@
with probability at least $1-\exp(-\zeta_{0}m\log(\frac{e}{\zeta_{0}}))$. Combining (3.21), (3.22) and (3.23), using small enough $\zeta_{0},c_{2}$, we obtain $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}\leq 5\sqrt{\zeta_{0}\log(e/\zeta_{0})}$.
Combining the two prior steps, we invoke Proposition 2.1 to obtain $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 37\sqrt{\zeta_{0}\log(e/{\zeta_{0}})}$. This leads to $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq 71\sqrt{\zeta_{0}\log(e/\zeta_{0})}$ by re-iterating (3.8). ∎
@@
Simulation: We pause to provide numerical evidence on the sharpness of $\tilde{O}(\sqrt{\zeta_{0}})$ for $\mathbf{x}^{\sharp}$, even under the optimally tuned noise level $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$. We adopt the same settings as in earlier simulations but replace the dense noise $\bm{\tau}$ by $\zeta_{0}m$ adversarial phase corruptions. We test $\zeta_{0}m=\{1,2,3,5,7,9,11,13\}$ and corrupt the measurements through the mechanism described in Proposition 3.1. The log-log curve in Figure 3(Left) roughly has a slope of $\frac{1}{2}$ over small $\zeta_{0}$. This seems to suggest the tightness of $\tilde{O}(\sqrt{\zeta_{0}})$ for the specific estimator $\mathbf{x}^{\sharp}$.
@@
As related context, in 1-bit compressed sensing, $\zeta_{0}m$ adversarial bit flips increment the $\ell_{2}$ error of the convex relaxation approach [40] by $\tilde{O}(\sqrt{\zeta_{0}})$, which was then improved to $\tilde{O}(\zeta_{0})$ using different algorithms [12, 11, 34, 1], and this (almost) linear increment is near-optimal under Gaussian designs (e.g., see [39, Thm. 2.4]).
@@
Consider $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$, fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and given $\delta\in(0,1)$. For some absolute constants $C_{1}$ and $c_{2}$, if $m\geq C_{1}\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)})$, then $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta)$ with probability at least $1-\exp(-c_{2}\delta^{2}m)$.
@@
We first show $\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\mathcal{U}_{c},\frac{1}{3})$ where $\mathcal{U}_{c}=\Sigma^{n}_{2s}\times\Sigma^{m}_{2\zeta_{0}m}$. By $\mathcal{U}_{c}^{\rm(N)}\subset(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})\times(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{n})$, we have $\omega(\mathcal{U}_{c}^{{\rm(N)}})\leq\omega(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+\omega(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})$ and hence
$\omega^{2}(\mathcal{U}_{c}^{\rm(N)})\leq 2\omega^{2}(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+2\omega^{2}(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})\leq C_{0}\Big{(}s\log\big{(}\frac{en}{s}\big{)}+\zeta_{0}\log\big{(}\frac{e}{\zeta_{0}}\big{)}m\Big{)}$
for some absolute constant $C_{0}$. Therefore, when $\zeta_{0}$ is sufficiently small, $m\geq C_{1}s\log(\frac{en}{s})$ with large enough $C_{1}$ implies $m\geq C_{2}\omega^{2}(\mathcal{U}_{c}^{\rm(N)})$ with large enough $C_{2}$. Then, by Theorem 3.7 we have $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},c_{3})$ with some small enough $c_{3}\leq\frac{1}{12}$ with probability at least $1-\exp(-c_{4}m)$. All that remains is to ensure $\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\rm(N)}}\|(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\|_{2}$ to be sufficiently small. Comparing $\mathbf{A}_{\mathbf{w}}$ in (1.7) and $\widetilde{\mathbf{A}}_{\mathbf{w}}$ in (3.26), we have
$\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\rm(N)}}\big{\|}(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\big{\|}_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\big{\|}(\mathbf{A}_{\breve{\mathbf{z}}}-\mathbf{A}_{\mathbf{z}})\mathbf{u}\big{\|}_{2}=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}.$
@@
with probability at least $1-2\exp(-c_{5}\zeta_{0}\log(\frac{e}{\zeta_{0}})m)$; see (3.10)–(3.12). Thus, $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ is small enough by the scaling of $m$ and small enough $\zeta_{0}.$
@@
where the last two inequalities hold with the promised probability due to the sub-Gaussian concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ and the proven sufficiently small $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$. ∎

### v3

@@ 3.2 Bounded Dense Noise
Thus, we obtain $|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|\leq 2m\tau_{0}$ and $\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}\leq 2\sqrt{m}\tau_{0}$, which, together with the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ in Equation (3.2), yield (3.21). The proof is complete. ∎
The next theorem provides converse bounds that indicate the sharpness of the $O(\tau_{0})$ bounds in Theorems 3.33.4. Using the complex Gaussian $\bm{\Phi}$ and observations $\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$ or $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\tau}$, we show that no algorithm can reconstruct $\mathbf{x}\in\Sigma^{n,*}_{s}$ to an error substantially smaller than $O(\tau_{0})$. The idea is to identify another signal $\mathbf{x}^{\prime}\in\Sigma^{n,*}_{s}$ that is indistinguishable from $\mathbf{x}$ and satisfies $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}=\tilde{\Omega}(\tau_{0})$.
@@
On the event $\{|\mathcal{J}_{\mathbf{x},s/4m}|\leq\frac{s}{4}\}$, (3.27) translates into (no more than) $\frac{s}{2}$ real linear equations, so such $\tilde{\bm{\delta}}$ exists when $s\geq 4$. Now we consider $\mathbf{x}^{\prime}=\frac{\mathbf{x}+\tilde{\bm{\delta}}}{\|\mathbf{x}+\tilde{\bm{\delta}}\|_{2}}\in\Sigma^{n,*}_{s},$ and similarly to (3.26) we have $\frac{\tau_{*}}{2}\leq\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}\leq\frac{3\tau_{*}}{2}$, and $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ holds with probability at least $1-\frac{4}{m}$. We further bound $\|\operatorname{sign}(\bm{\Phi}\mathbf{x}^{\prime})-\operatorname{sign}(\bm{\Phi}\mathbf{x})\|_{\infty}$ by discussing two cases:
  • For $i\in\mathcal{J}_{\mathbf{x},s/4m}$, we have $\bm{\Phi}_{i}^{*}\tilde{\bm{\delta}}=0$ and hence $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}^{\prime})=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$.
  • For $i\notin\mathcal{J}_{\mathbf{x},s/4m}$, (2.11) gives
@@
    where the last inequality we use $\|\bm{\Phi}(\mathbf{x}^{\prime}-\mathbf{x})\|_{\infty}\leq 6\sqrt{\log m}\cdot\tau_{*}$ and substitute $m=C_{0}s\log(\frac{en}{s})$.
Taken collectively, it follows that
@@
Simulation: [^7] The MATLAB code used to generate the figures in this paper is available at https://github.com/junrenchen58/instance-optimal-pocs. We pause to use experimental results to provide evidence of the achievability and tightness of $O(\tau_{0})$. In all of our experiments, the data points are averaged over $50$ independent trials, each of which concerns the recovery of $\mathbf{x}$ uniformly drawn from $\Sigma^{500,*}_{5}$ using $300$ phase-only measurements. We provide the optimally tuned $\varepsilon$ to basis pursuit (1.8), namely $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$. For the post-sign noise, we test $\tau_{0}\in\{0.04,0.08,0.12,0.16,\cdots,0.36,0.40\}$ and adopt such corruption pattern: find $\theta_{0}\in[0,\frac{\pi}{2}]$ such that $|e^{\mathbf{i}\theta_{0}}-1|=\tau_{0}$ and then corrupt $\mathbf{z}$ to $\breve{\mathbf{z}}=e^{\mathbf{i}\theta_{0}}\mathbf{z}$. For the pre-sign noise, we test $\tau_{0}\in\{0.04,0.12,0.20,0.28,\cdots,0.76,0.84\}$ and generate the noisy observations through $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\tau_{0}\operatorname{sign}(\bm{\Phi}\mathbf{x})\mathbf{i})$. The results are given in Figures 12 and are consistent with our theorems.
@@
Remark 3.4 (The choice of noises in simulations).
Note that $O(\tau_{0})$ in Theorems 3.3, 3.4 is a worst-case upper bound that applies to arbitrary $\bm{\tau}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$, but it may not be tight for some $\bm{\tau}$: for instance, under the pre-sign noise $\bm{\tau}=\tau_{0}\operatorname{sign}(\bm{\Phi}\mathbf{x})$, the observations $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ are identical to the noiseless ones $\mathbf{z}$, and therefore perfect recovery is achieved and $O(\tau_{0})$ is not tight. In view of Lemma 3.1, the final error is at the same order of $\varepsilon$ obeying $\varepsilon\geq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$. To numerically observe the tightness of $O(\tau_{0})$, the noises in Figures 12 are chosen such that $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\Omega(\tau_{0})$ with high probability.
@@ 3.3 Sparse Phase Corruption
Consider PO-CS of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\bm{\zeta}$ with $\bm{\zeta}$ obeying $\|\bm{\zeta}\|_{\infty}\leq 2$ and $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ for some $\zeta_{0}\leq c_{0}$ where $c_{0}$ is a small enough absolute constant. If $m\geq C_{1}s\log(\frac{en}{s})$ with sufficiently large $C_{1}$, then the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$, with $\hat{\mathbf{x}}$ being solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})})$ in (1.8), satisfies
@@
By Lemma 3.2, it suffices to establish (3.7) and (3.3) with $\varepsilon=5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})}$, then we have (3.8) that reads $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}\cdot 70\sqrt{\zeta_{0}\log(e/\zeta_{0})}}{\kappa m}$. This yields the claim due to $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}<\frac{71}{70}$ that holds with the promised probability (see Equation (3.2)).
Establish (3.7): In the proof of Theorem 3.4 we show (3.20) for $\tilde{\bm{\tau}}_{2}$ satisfying $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$ and $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$. For $\bm{\zeta}=\breve{\mathbf{z}}-\mathbf{z}$ obeying $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$ in the current setting, an identical argument yields
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\zeta_{0}\log(e/\zeta_{0})}\right)$
with the probability we promise. This yields (3.7) by the scaling of $m$ and $\zeta_{0}\leq c_{0}$ for some small enough $c_{0}$.
Establish (3.3) with $\varepsilon=5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})}$: We let $\mathbf{1}_{\operatorname{supp}(\bm{\zeta})}\in\{-1,1\}^{m}$ whose $1$’s indicate the support set of $\bm{\zeta}$. Then we have
@@
Without loss of generality, we assume $\zeta_{0}m$ is a positive integer. For fixed $I$ with cardinality $\zeta_{0}m$, $\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}$ follows Chi-squared distribution with $2\zeta_{0}m$ degrees of freedom. Then by a standard concentration bound [31, Lem. 1], we obtain that for any $t\geq 0$ and any $I\subset[m]$ with $|I|=\zeta_{0}m$,
@@
with probability at least $1-\exp(-\zeta_{0}m\log(\frac{e}{\zeta_{0}}))$. Substituting (3.30) into (3.29), using small enough $\zeta_{0}$ and $\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\leq 1+c$ for small enough $c$ due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ in Equation (3.2), we establish (3.3) with $\varepsilon=5\sqrt{\zeta_{0}\log(\frac{e}{\zeta_{0}})}$. ∎
@@
Simulation: We pause to provide numerical evidence on the sharpness of $\tilde{O}(\sqrt{\zeta_{0}})$ for $\mathbf{x}^{\sharp}$, even under the optimally tuned noise level $\varepsilon=\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$. We adopt the same settings as in earlier simulations (i.e., $m=300$, $n=500$, $s=5$) but replace the dense noise $\bm{\tau}$ by $\zeta_{0}m$ adversarial phase corruptions. We test $\zeta_{0}m\in\{1,2,3,5,7,9,11,13\}$ and corrupt the measurements through the mechanism described in Proposition 3.1. The log-log curve in Figure 3(Left) roughly has a slope of $\frac{1}{2}$ over small $\zeta_{0}$. This seems to suggest the tightness of $\tilde{O}(\sqrt{\zeta_{0}})$ for the specific estimator $\mathbf{x}^{\sharp}$.
@@
As related context, in 1-bit compressed sensing, $\zeta_{0}m$ adversarial bit flips increment the $\ell_{2}$ error of the convex relaxation approach [42] by $\tilde{O}(\sqrt{\zeta_{0}})$, which was then improved to $\tilde{O}(\zeta_{0})$ using different algorithms [12, 11, 35, 1], and this (almost) linear increment is near-optimal under Gaussian designs (e.g., see [40, Thm. 2.4]).
@@
Consider $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$, fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and given $\delta\in(0,1)$. For some absolute constants $C_{1}$ and $c_{2}$, if $m\geq C_{1}\delta^{-2}\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})$, then $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},\delta)$ with probability at least $1-\exp(-c_{2}\delta^{2}m)$.
@@
We first show $\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\mathcal{U}_{c},\frac{1}{3})$ where $\mathcal{U}_{c}=\Sigma^{n}_{2s}\times\Sigma^{m}_{2\zeta_{0}m}$. By $\mathcal{U}_{c}^{\mathbb{S}}\subset(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})\times(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{n})$, we have $\omega(\mathcal{U}_{c}^{\mathbb{S}})\leq\omega(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+\omega(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})$ and hence
$\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})\leq 2\omega^{2}(\Sigma^{n}_{2s}\cap\mathbb{B}_{2}^{n})+2\omega^{2}(\Sigma^{m}_{2\zeta_{0}m}\cap\mathbb{B}_{2}^{m})\leq C_{0}\Big(s\log\big(\frac{en}{s}\big)+\zeta_{0}\log\big(\frac{e}{\zeta_{0}}\big)m\Big)$
for some absolute constant $C_{0}$. Therefore, when $\zeta_{0}$ is sufficiently small, $m\geq C_{1}s\log(\frac{en}{s})$ with large enough $C_{1}$ implies $m\geq C_{2}\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})$ with large enough $C_{2}$. Then, by Theorem 3.7 we have $\widetilde{\mathbf{A}}_{\mathbf{z}}\sim\operatorname{RIP}(\mathcal{U}_{c},c_{3})$ with some small enough $c_{3}\leq\frac{1}{12}$ with probability at least $1-\exp(-c_{4}m)$. All that remains is to ensure $\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\mathbb{S}}}\|(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\|_{2}$ to be sufficiently small. Comparing $\mathbf{A}_{\mathbf{w}}$ in (1.7) and $\widetilde{\mathbf{A}}_{\mathbf{w}}$ in (3.33), we have
$\sup_{\mathbf{u}\in\mathcal{U}_{c}^{\mathbb{S}}}\big\|(\widetilde{\mathbf{A}}_{\breve{\mathbf{z}}}-\widetilde{\mathbf{A}}_{\mathbf{z}})\mathbf{u}\big\|_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\big\|(\mathbf{A}_{\breve{\mathbf{z}}}-\mathbf{A}_{\mathbf{z}})\mathbf{u}\big\|_{2}=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}.$
@@
with probability at least $1-2\exp(-c_{5}\zeta_{0}\log(\frac{e}{\zeta_{0}})m)$; see Equations (3.18)–(3.20). Thus, $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ is small enough by the scaling of $m$ and small enough $\zeta_{0}.$
@@
where the last two inequalities hold with the promised probability due to the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$ in Equation (3.2) and the sufficiently small $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$. ∎